OpenStudy (1018):

(physics) You drive on Interstate 10 from San Antonio to Houston, half the time at 79 km/h and the other half at 105 km/h. On the way back you travel half the distance at 79 km/h and the other half at 105 km/h. What is your average speed (a) from San Antonio to Houston, (b) from Houston back to San Antonio, and (c) for the entire trip? (d) What is your average velocity for the entire trip?

OpenStudy (1018):

my answer for (a) is (79+105)/2 = 92 and (d)= 0. are these correct?

OpenStudy (anonymous):

Yes.

OpenStudy (1018):

thanks. but how do i solve for b and c?

OpenStudy (anonymous):

b) is done exactly the same way as a). What do you think?

OpenStudy (1018):

yeah, that's what i thought. that it would be tha same but it was wrong when i tried it.

OpenStudy (1018):

i also found a similar question with different values, and they a and b werent the same either

OpenStudy (anonymous):

Can't understand why. Half the trip is at 79 and the other half is at 105.

OpenStudy (1018):

i know! me neither

OpenStudy (1018):

and c would be just a+b / 2 it was also wrong

OpenStudy (anonymous):

Oh, I get it now. The difference is, for the return trip, half the DISTANCE is at 79 and the other at 105. Gotta read carefully.

OpenStudy (1018):

OOOH I didnt see that hahaha

OpenStudy (1018):

wait ill try haha. can you answer it too? just to see if im right

OpenStudy (anonymous):

Sure. Go ahead.

OpenStudy (anonymous):

for 2nd consider the the distance between the 2 places be \[s=2x\] Assume time taken \[t_{1}\] for first half of the distance and time \[t_{2}\] for the 2nd half \[v_{av}=\frac{s}{t}=\frac{s}{t_{1}+t_{2}}\] Let us find t1 and t2, \[t_{1}=\frac{s/2}{79}=\frac{x}{79}\] similarly \[t_{2}=\frac{x}{105}\]\[v_{av}=\frac{2x}{\frac{x}{79}+\frac{x}{105}}\] On simplification the x will be cancelled, alternatively you could do this with just the s and without introducing the x

OpenStudy (1018):

i got 90? is that correct? @Nishant_Garg

OpenStudy (anonymous):

that's a good enough answer, it will not come as a round figure, I'm getting around 89.19

OpenStudy (1018):

is it ok? originally i got 90.12 something. was that wrong or that's acceptable already?

OpenStudy (1018):

oh nevermind, 90 was correct. haha. thanks!

OpenStudy (anonymous):

for 3rd I assume you take the average of the forward and back trip's speed's

OpenStudy (1018):

yes that's what i did, and it's correct

OpenStudy (anonymous):

actually yeah, I made a slight mistake 90.16 is the correct answer so approximately 90

OpenStudy (anonymous):

for the 2nd part

OpenStudy (anonymous):

so for 3rd it should average of 90 and 92, so 91

OpenStudy (1018):

yes, all answers were right. thanks!